OpenStudy (anonymous):

A square piece of fabric has an area of 595 cm2. To the nearest tenth, what is the length of 1 side of the fabric? A. 21.3 cm B. 23.2 cm C. 24.4 cm D. 25.6 cm

OpenStudy (campbell_st):

just take the square root of the area. That will give the side length remember is a square side length l \[Area = l^2\] so if you know the area \[595 = l^2\] hope it helps

OpenStudy (anonymous):

okay so do i divide 595 by the sq root?

OpenStudy (anonymous):

@campbell_st ?

OpenStudy (campbell_st):

no you take the square root... as an example \[\sqrt{16} = 4 \]

OpenStudy (anonymous):

oh okay thats what i ment THANKS

OpenStudy (anonymous):

c?

OpenStudy (campbell_st):

I think so... 25^2 = 625... so its between 24 and 25

OpenStudy (anonymous):

wait if there are four sides would i have to separate 595 into four

OpenStudy (campbell_st):

no... in a square every side is the same length... so area is Area = length x length or A = l^2 so take the square root of the area to get the side length you answer is correct

OpenStudy (anonymous):

thanks for helping me understand XD